OpenStudy (anonymous):

integrate it:dx/(2cosec.x-sin.x)

OpenStudy (anonymous):

\[\int\limits_{?}^{?} [1/ (cosec(x) -\sin(x) ]dx\]

OpenStudy (anonymous):

is this the correct equation /?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first convert the cosec to 1 / cosx

OpenStudy (anonymous):

sorry 1/ sin x

OpenStudy (anonymous):

\[dx \div(2coces.x-\sin.x)\]

OpenStudy (anonymous):

then multiply so you will get (1 - sin^2 x)/Sin x

OpenStudy (anonymous):

then

OpenStudy (anonymous):

then 1 - sin^2 x = co^2 x

OpenStudy (anonymous):

so you will end up with cos^2x /sinx

OpenStudy (anonymous):

the end integration will be some thing like \[\int\limits_{?}^{?} ( \cos^2x ) / \sin x dx\]

OpenStudy (anonymous):

which can solved easily after that